Home › CBSE Class 12 (Commerce) › Mathematics › Probability › E1, E2 partition S: P(E1)=0.6, P(A|E1)=0.2, P(E2…
E1, E2 partition S: P(E1)=0.6, P(A|E1)=0.2, P(E2)=0.4, P(A|E2)=0.5. Find P(A).
A0.7
B0.26
C0.32
D0.40
Answer & Solution
Correct answer: C. 0.32
Total probability gives P(A) = 0.6 x 0.2 + 0.4 x 0.5 = 0.12 + 0.20 = 0.32.
Related questions
If P(A) = 0.6 and P(B) = 0.7, the smallest possible P(A and B) isA bag has 5 red, 3 green and 2 yellow marbles. P(red or yellow) isTwo dice are thrown. P(first die is 6 and sum is at least 11) isA card is drawn from ten cards numbered 1 to 10. P(2, 4, 6 or 8) isTwo fair dice are thrown. P(sum greater than 11) isTwo fair coins are tossed. P(exactly one head) isA fair die is rolled. P(a number less than 3) isIn P(A) + P(B), the sample points of A and B together are counted